OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–8, 3); Slope: 6 A.y + 3 = 6x – 48 B.y + 3 = 6(x + 8) C.y – 3 = 6(x – 8) D.y – 3 = 6(x + 8) And Which is an equation in point-slope form for the given point and slope? Point: (2, -6) ; Slope: -3/4 A. y - 6 = -3/4(x+2) B. y - 6 = 3/4(x - 2) C. y + 6 = -3/4(x - 2) D. y + 6 = -3/4(x + 2) As usual anyone who answers I will mark them best post & give them a +fan. All help is very appreciated. Thanks! :)

OpenStudy (anonymous):

Use the formula: \[\frac{y-y1}{x-x1} = m\] (where m = slope) \[\frac{y-3}{x-(-8)} = 6\]\[\frac{y-3}{x-(-8)} *( x - (-8)) = 6* (x - (-8))\] \[D. y-3 = 6(x + 8)\] \[\frac{y-(-6)}{x-2} = -\frac{3}{4}\] \[C. y+6 = -\frac{3}{4} * (x-2)\]

OpenStudy (anonymous):